TEA Banner
TEA Navbar

3 December, 2002

Composite of science today

Date: 12/03/02

Latitude: 80 degrees South

Longitude: 120 degrees West

Temperature: -14 C / +7 F

Wind speed: 12 knots / 14 MPH

Wind Chill: - 23=B0C / --10=B0F

Wind direction: North Northeast

Visibility: forever!

Conditions: Clear and sunny

Meters of ice collected so far: 37.5 m

Notes on daily life by Betsy and Dan

Today in the atmospheric chemistry section of the camp, situated one and a half kilometers NW of Byrd camp, Markus and Betsy have completed their set up and testing of their suite of instruments and experiments. Late last night Markus completed the set up of his weather observing instruments. These include a radiometer to collect UV data, an anemometer to sense wind speed and direction, and the remainder of his atmospheric chemistry detectors. The winds calmed enough this morning so that it was possible to try out the tethered balloon launch again. Several successful tethered launches were achieved as well as one unanticipated free (untethered) launch.

The balloon had been stored in a snow pit for the past several days while they waited for the winds to die down below 15 knots. After lunch Markus and Betsy hiked back out to the chemistry zone and while Markus calibrated his instruments, Betsy went to work to prepare for a launch. Brian skied out a little while later to help out with digging another snow pit and managing the balloon. The balloon had been leaking helium over the past few days and it would not lift the sonde and O3 sensing equipment. They reeled it back in to add more helium, but the sunlight had weakened the latex balloon sufficiently so that it popped from the pressure. Their second balloon had a small hole in it which kept it from getting to ideal altitude, they reeled it back in and tied another balloon to it. Unfortunately, the double balloon overpowered Betsy's knot and the balloon pulled itself free from the wooden plug that keeps the helium from escaping the balloon while in flight. After this lighter moment, they all refocused and sent off one last tethered flight allowing them to finally gather the desired data about the atmospheric conditions (temperature, relative humidity, barometric pressure, ozone concentrations) several hundred feet above the snow.

Meanwhile the drill team was having its own adventures. Mark's revolutionary 2-inch drill set a new depth record 37.5m. The ice core from this depth will contain a paleoclimate record extending back in time over 200 years. Also this morning, Blue completed the last 100 km of his snowmobile GPS survey and in the afternoon he and Gordon set up a strain grid around the camp.

Andrea prepared another fantastic feast for supper and we are now watching "Kingpin" on DVD.

Contact the TEA in the field at .
If you cannot connect through your browser, copy the TEA's e-mail address in the "To:" line of your favorite e-mail package.